OpenStudy (mokeira):

If the roots of the equation x^2+bx+c=0 are two consecutive integers, what is the value of b^2-4c-1

hartnn (hartnn):

hi do you know sum of roots for \(ax^2+bx+c=0\) is -b/a and product of roots =c/a ?

OpenStudy (mokeira):

no..i didnt! Thanks for that. let me try now and see if I will get stuck

hartnn (hartnn):

you will also need this : \((p-q)^2 = (p+q)^2 -4pq\) p and q being the roots of the equation

hartnn (hartnn):

and since roots are consecutive p-q =\(\pm 1\)

OpenStudy (mokeira):

umm...can you just take me through step by step

hartnn (hartnn):

in this case a=1 right ?

hartnn (hartnn):

so, p+q = -b pq = c makes sense till now ?

OpenStudy (ikram002p):

lot of sense !

OpenStudy (mokeira):

what does p+q represent

hartnn (hartnn):

p and q are the roots of the equation so p+q = sum of the roots = -b/a =-b

OpenStudy (mokeira):

ok...i get it now

hartnn (hartnn):

can you try to continue ? use the formula i gave above p-q =1 pq =c p+q =-b

OpenStudy (mokeira):

let me show you how I have understood and interpreted. if the roots are (m) and (m+1), where m is an integer then: sum of roots= (m)+(m+1)=2m+1=-b product of roots= (m)(m+1)= \[m ^{2}+m=c \] so \[b ^{2}=(-b)^{2}=[-(2m+1)]^{2}= 4m ^{2}+4m+1\] Therefore \[b ^{2}-4c-1=(4m ^{2}+4m+1)-4(m ^{2}+m)-1\] am i on the right track?

hartnn (hartnn):

absolutely! you just did it the other way , but you will get the correct answer after you simplify :)

OpenStudy (mokeira):

yaaaaay! thanks

hartnn (hartnn):

welcome ^_^